OpenStudy (anonymous):

How many pounds of candy selling for $1.75 a pound should be added to three pounds of candy selling for $4.50 a pound to create a blend, which will sell for $3.40 a pound?

OpenStudy (mathmate):

Mixing/blending problems. To solve these problems, set up an equation relating the total price of all the candies, and divide by the total weight. Represent the unknown by a variable (such as x) and solve for x. "How many pounds of candy selling for $1.75 a pound" So call this "how many" x (in pounds), so Total cost of cheaper candy = 1.75x total cost of expensive candy = 4.50(3) Total cost = 1.75x+3(4.50) Total weight = 3+x Unit price of mixture = total cost / total weight = (1.75x+3(4.50)) / (3+x) = 3.40 So solve for x Cross multiply: (1.75x+13.50)=3.40\(\times\)(3+x) I will leave it to you to solve for x, which is the weight of candy costing $1.75 a pound. Note: as a rough check, the unit price of the mixture must be between unit prices of the two components. If not, there is probably a mistake.

OpenStudy (anonymous):

@mathmate what did you end up getting for x?

OpenStudy (mathmate):

Post your answer and I'll be glad to check it for you. Please also include how you solved it.